• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Moffit

    Log entry time 03:16:56 on November 10, 2008

    Entry number 247896

    This entry is a followup to: 247891


    keyword=Re: Feedback not working... Beam Charge asymmetry still small.

    I just checked the current run 30393. Even though feedback is not working... the beam charge asymmetry appears to be small.

    I see less than 50 ppm from the downstream x3 in the HAPPEX DAQ... there's a discrepency with the HRS DAQ numbers, but I think that this is because the lack of beam cuts used in the End-of-Run entries.